# Accelerating Multimodal Large Language Models via Dynamic Visual-Token Exit and the Empirical Findings

### v1

@@ 4.4 Qualitative Analysis
To gain insight into the proposed DyVTE, we visualize its predictions and inference efficiency on LLaVA-1.5 7B in Fig.7. We can first observe that the implementation of DyVTE maintains consistency with the default LLaVA. As shown in Fig.7-(a), DyVTE can correctly identify the object in question, even with a complex foreground. Additionally, it is clear that DyVTE provides much faster inference than the default LLaVA. For two examples, DyVTE can improve the inference speed by up to 10.2% compared to the default LLaVA. Another observation is that removing redundant visual information can further enhance the LLaVA’s performance. As shown in Fig.7-(b), similar to the results on the MM-Vet benchmark, removing visual tokens also improves the performance of the default LLaVA. Overall, these results confirm that DyVTE enhances inference efficiency without compromising the reasoning process of the default model, aligning well with our motivation.
@@ 5 Conclusion
In this paper, we investigate the attention patterns in multimodal large language models (MLLMs) and propose a a simple yet effective method, dynamic visual-token exit (DyVTE), to improve computational efficiency. By analyzing the attention patterns in the MLLMs, we identify three key stages of most MLLMs in the attention process, i.e., early fusion, inter-modality modeling, and multimodal reasoning. Based on this finding, we further show show that visual tokens contribute less to the reasoning process as inference progresses. Motivated by this insight, DyVTE optimally determines the layer at which visual tokens can be removed, thereby reducing computational overhead without compromising performance. Extensive experiments across five MLLMs and VL benchmark benchmarks demonstrate that DyVTE significantly enhances the efficiency of MLLMs while maintaining their performance.
6 Acknowledgments
This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001).
@@ Appendix A Appendix
A.1 Ablation Study
In Tab.5 and Tab.6, we perform a set experiment to analyze the impact of different representation selections on DyVTE. As shown in Tab.5, when relying on more than two representations, we can find out that their performance is similar to “mean text + last text”. For instance, the additional representation, i.e., “last visual” and “mean visual”, has no noticeable effect on the results. We can also observe that the missing of “mean text” or “last text” will lead to inaccurate exit. For example, the missing of “mean text” leads to later exit, i.e., 2 layers latter, while the absence of “last text” does the opposite, i.e., 2 layers earlier. As for taking attention scores be the representations, we find that hyper-network can not make the correct judgment. Specifically, the exit layers for all benchmarks can be much earlier. Overall, these results well confirm the use of text tokens status for dynamic visual token exiting in MLLMs.
@@ A.1 Ablation Study
Mean Visual Last Visual Mean Text Last Text Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$
@@
Table 6: Ablation study of the attention status selection for DyVTE. “Visual Self” denotes the attention score from visual modality of each visual tokens, similar with “Text Self”. “Cross” denotes the attention score from visual to the text of each visual tokens. For attention representations whose dimensions are not 576, we use interpolation to align the dimensions. “Exit Layer” denotes the averaged layer numbers selected to remove by DyVTE. “Acc.” denotes the accuracy. The default setting of DyVTE is the last row. The best and second best results are marked in bold and underline, receptively.
@@
Visual Self Cross Text Self Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$ Acc. $\uparrow$ Exit Layer $\downarrow$
@@
A.2 Distribution of Attention
@@
A.3 Entropy of Attention
In Fig.13-16, we visualize the entropy distributions of cross-attention matrices and text self-attention matrices. For all these MLLMs and datasets, we can observe that removing all visual tokens at an appropriate time will not have a significant impact on the answer modeling process. Specifically, for all MLLMs and datasets, the removing of visual tokens only makes the entropy of cross-attention disappear, while the entropy of text self-attention maintains the same distribution. Overall, these results well confirm that cross-modal interaction in fact barely contributes to multimodal reasoning after some layers.

### v2

@@ 5.4 Qualitative Analysis
Figure 6: Examples on LLaVA-1.5 with DyVTE. Our DyVTE can help the MLLM answer the questions as accurately as default, while being faster.
To gain insight into the proposed DyVTE, we visualize the effect of applying it on LLaVA-1.5 7B in Fig. 6. We can first observe that DyVTE maintains consistent predictions with the default LLaVA. As shown in Figure 6-(a), DyVTE can correctly identify the object in question even when the foreground is complex. Additionally, DyVTE provides much faster inference than the default LLaVA. Specifically, in the two given examples, DyVTE can improve inference speed by up to 10.2% compared to the default LLaVA. Another observation is that removing redundant visual information may enhance performance. As shown in Fig.6-(b), applying DyVTE can help LLaVA to better recognize the numbers on the sign. Overall, these results confirm that DyVTE enhances inference efficiency without harming the reasoning process of the default model, aligning well with our motivation.
6 Limitation
While DyVTE shows promising efficiency gains across multiple MLLMs and benchmarks, DyVTE currently performs full visual-token exit, which may not be optimal for cases where partial visual information is still beneficial in later layers.
@@ 7 Conclusion
In this paper, we investigate the visual redundancy problem of MLLMs via analyzing their inference behaviors, and summarize three key stages of MLLMs during multimodal inference, i.e., early fusion, inter-modality modeling, and multimodal reasoning. Moreover, we also reveal that visual tokens stop contributing to multimodal reasoning after some layers. Motivated by this insight, we propose a simple yet effective method for MLLMs, termed Dynamic Visual-Token Exit (DyVTE) which optimally determines the layer at which visual tokens can be removed, thereby reducing computational overhead without compromising performance. Extensive experiments on nine benchmarks demonstrate that DyVTE significantly enhances the efficiency of MLLMs while maintaining their performance.
@@ Appendix A Appendix
A.1 Attention Visualization
In Fig.2, we visualize the attention weight of each part in attention matrix, and the attention weights are calculated as follow: We first calculate the attention matrix with system prompt $\mathbf{S}\in\mathcal{R}^{s\times d}$, visual tokens $\mathbf{V}\in\mathcal{R}^{v\times d}$ and text tokens $\mathbf{T}\in\mathcal{R}^{t\times d}$:
$\mathbf{A}=Softmax(\frac{[\mathbf{S},\mathbf{V},\mathbf{T}]\mathbf{W}_{q}([\mathbf{S},\mathbf{V},\mathbf{T}]\mathbf{W}_{k})^{T}}{\sqrt{d}}\cdot\mathbf{M}),$ (7)
where $\mathbf{M}$ represents the attention mask for causal reasoning. Then, the weight of each part can be represent as
$W_{v}=\sum_{i=s}^{s+v}\sum_{j=s}^{s+v}\mathbf{A}_{ij}/v,W_{c}=\sum_{i=s+v}^{s+v+t}\sum_{j=s}^{s+v}\mathbf{A}_{ij}/t,W_{t}=\sum_{i=s+v}^{s+v+t}\sum_{j=s+v}^{s+v+t}\mathbf{A}_{ij}/t$ (8)
where $W_{v}$, $W_{c}$ and $W_{t}$ represent the attention weights of visual self attention, cross attention and text self attention, respectively. In addition, we combine the attention matrices of different heads in MHA through the max pooling.
A.2 Ablation Study
In Tab.6 and Tab.7, we perform a set of experiments to analyze the impact of different representation selections on DyVTE. As shown in Tab.6, when relying on more than two representations, we can find out that their performance is similar to “mean text + last text”. For instance, the additional representation, i.e., “last visual” and “mean visual”, has no noticeable effect on the results. We can also observe that the missing of “mean text” or “last text” will lead to inaccurate exit. For example, the missing of “mean text” leads to later exit, i.e., 2 layers latter, while the absence of “last text” does the opposite, i.e., 2 layers earlier. As for taking attention scores be the representations, we find that hyper-network can not make the correct judgment. Specifically, the exit layers for all benchmarks can be much earlier. Overall, these results well confirm the use of text tokens status for dynamic visual token exiting in MLLMs.
@@ A.2 Ablation Study
Mean
Visual
Last
Visual
Mean
Text
Last
Text
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
@@
Table 7: Ablation study of the attention status selection for DyVTE. “Visual Self” denotes the attention score from visual modality of each visual tokens, similar with “Text Self”. “Cross” denotes the attention score from visual to the text of each visual tokens. For attention representations whose dimensions are not 576, we use interpolation to align the dimensions. “Exit Layer” denotes the averaged layer numbers selected to remove by DyVTE. “Acc.” denotes the accuracy. The default setting of DyVTE is the last row. The best and second best results are marked in bold and underline.
@@
Visual
Self
Cross
Text
Self
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
Acc. $\uparrow$
Exit
Layer
$\downarrow$
@@
A.3 Distribution of exit layers
In Fig. 7, we present the distributions of the visual token exit layers for LLaVA-DyVTE 7B and 13B. From the figure, we observe that DyVTE dynamically selects the exit layer based on the specific task requirements. For example, on the TextVQA dataset, which demands fine-grained information, the exit occurs later compared to other datasets. In contrast, on the GQA dataset, which involves open-ended word questions and simple true/false judgments, the exit layers are more evenly distributed across early and late layers. Another notable observation is that the timing of the exit layer shows similar distributions across MLLMs of different scales and architectures. For instance, both LLaVA-DyVTE 7B and 13B remove all visual tokens at the same layer early in the process on the SQA dataset. Similarly, for InternVL and LLaVA, peaks in the exit layer distributions occur at comparable layers. Overall, these results support our motivation and the proposed DyVTE.
Figure 7: Statistics of the exiting layers decided by DyVTE on LLaVA-1.5 7B, LLaVA-1.5 13B and InternVL 7B. The horizontal axis represents the exit layer, and vertical axis represents the proportion. The exit time by is different for different tasks, but similar for MLLMs with the same sizes.
A.4 Distribution of Attention
@@
A.5 Entropy of Attention
In Fig.14-17, we visualize the entropy distributions of cross-attention matrices and text self-attention matrices. For all these MLLMs and datasets, we can observe that removing all visual tokens at an appropriate time will not have a significant impact on the answer modeling process. Specifically, for all MLLMs and datasets, the removing of visual tokens only makes the entropy of cross-attention disappear, while the entropy of text self-attention maintains the same distribution. Overall, these results well confirm that cross-modal interaction in fact barely contributes to multimodal reasoning after some layers.
@@ A.5 Entropy of Attention
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction explicitly outline the main contributions, including an empirical findings in the inference MLLM and a novel and effective DyVTE method to exit the visual tokens in an early stage.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation have been provided in Sec. 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include any new theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all the details in our paper, and our code is anonymously released.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code is anonymously released and all datasets we used is public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all the details in our paper, and our code is anonymously released.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The paper does not include experiments involving random processes that would require the reporting of error bars or statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provide these information in our experiment details.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have strictly adhered to the guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper does not explicitly discuss the societal impacts of the work performed, as it primarily focuses on the technical development of multimodal large language models.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work is based on public data and models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited their works.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our code is produced in our anonymous project.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [No]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: LLM, as a part of MLLM, is the object of our study.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.